The density of water at 4.00°C is 0.967 g/mL. How many molecules of water are present in a 499.8 mL bottle of water? Express your answer to the correct number of significant figures. Be sure to show all steps completed to arrive at the answer.



1
Expert's answer
2021-03-04T06:34:37-0500

Density of water at 4.00oC is 0.967 g/mL . Volume of water = 499.8 mL

Mass of water ="density \times volume = 0.967 \times 499.8=483.3066 g"

Number of moles of water molecules in given mass of water ="\frac{given mass}{molar mass}= \frac{483.3066}{18}=26.85"

we know that 1 mole has "6.022\times 10^{23}" molecules of water

so,

number of water molecules="6.022\times 10^{23} \times 26.85=161.69\times 10^{23}=1.62 \times 10^{25}" molecules of water. It is upto 3 significant figure.